# Feature-aware Hypergraph Generation via Next-Scale Prediction

### v1

@@ B.2 Minibatch OT-coupling
$f(\mathbf{x}_{0})=f(\sigma(\mathbf{x}_{0})).$
@@
Appendix C Complexity Analysis
[…]
Appendix D Coarsening Sequence Sampling
[…]
@@ Appendix D Coarsening Sequence Sampling
i.e., when the number of nodes on the left side (corresponding to the original hypergraph nodes) has been reduced by the sampled fraction. This framework is flexible, allowing a variety of cost functions $c$, contraction families $F$, reduction fraction ranges $[\rho_{\text{min}},\rho_{\text{max}}]$, and randomization parameters $\lambda$.
@@
  • Reduction fraction range: $[\rho_{\text{min}},\rho_{\text{max}}]=[0.1,0.3]$.
@@
1:contraction family $F$, cost function $c$, reduction fraction range $[\rho_{\text{min}},\rho_{\text{max}}]$, randomization parameter $\lambda$
@@
8:     $\mathbf{b}_{L}^{(0)}\leftarrow(1,\dots,1)\in\mathbb{R}^{n}$triangleright Initial node budgets
9:     $\mathbf{b}_{R}^{(0)}\leftarrow(1,\dots,1)\in\mathbb{R}^{m}$triangleright Initial hyperedge budgets
@@
[…]
15:         $f\leftarrow c(\cdot,C^{(l-1)},(\mathcal{P}^{(l-1)},\ldots,\mathcal{P}^{(0)}))$triangleright Cost function
[…]
@@
[…]
24:                           $C^{(l)}\leftarrow C_{\text{temp}}$, $B^{(l)}\leftarrow B_{\text{temp}}$
25$\triangleright$ Update budgets and features for the new cluster
26:                           Let $\text{Let }S=\{v_{1},\ldots,v_{k}\},\text{ and the new node be }v^{*}$
@@
Appendix E Training and Sampling Procedures
1:number of spectral features $k$
4:function Embeddings($B=(\mathcal{V}_{L},\mathcal{V}_{R},\mathcal{E})$, SignNetθ, $\mathbf{v}_{L}$, $\mathbf{v}_{R}$)
5:     if $k=0$ then
6:         $\mathbf{H}=[h^{(1)},\ldots,h^{(|\mathcal{V}|)}]\overset{i.i.d.}{\sim}\mathcal{N}(0,I)$triangleright Sample random embeddings
7:     else
8:         if $k<|\mathcal{V}|$ then
9:              $[\lambda_{1},\ldots,\lambda_{k}],[u_{1},\ldots,u_{k}]\leftarrow\text{EIG}(B)$triangleright Compute $k$ spectral features
10:         else
11:              $[\lambda_{1},\ldots,\lambda_{|\mathcal{V}_{L}|+|\mathcal{V}_{R}|-1}],[u_{1},\ldots,u_{|\mathcal{V}_{L}|+|\mathcal{V}_{R}|-1}]\leftarrow\text{EIG}(B)$triangleright Compute $|\mathcal{V}_{L}|+|\mathcal{V}_{R}|-1$ spectral features
12:              $[\lambda_{|\mathcal{V}_{L}|+|\mathcal{V}_{R}|},\ldots,\lambda_{k}],[u_{|\mathcal{V}_{L}|+|\mathcal{V}_{R}|},\ldots,u_{k}]\leftarrow[0,\ldots,0],[0,\ldots,0]$triangleright Pad with zeros
13:         end if
15:     end if
16:     $\tilde{B}=(\mathcal{V}_{L}^{(1)}\cup\cdots\cup\mathcal{V}_{L}^{(p_{l})},\mathcal{V}_{R}^{(1)}\cup\cdots\cup\mathcal{V}_{R}^{(p_{r})},\tilde{\mathcal{E}})\leftarrow\tilde{B}(B,\mathbf{v}_{L},\mathbf{v}_{R})$triangleright Expand as per Definition 4
17:     set $\tilde{B}$ s.t. for all $p_{L}\in[|\mathcal{V}_{L}|]$ and all $p_{R}\in[|\mathcal{V}_{R}|]$: for all $\mathbf{v}_{L}^{(p_{i})}\in\mathcal{V}_{L}^{(p_{l})}$, $\tilde{\mathbf{H}}[p_{i}]=\mathbf{H}[p_{l}]$ and for all $\mathbf{v}_{R}^{(p_{i})}\in\mathcal{V}_{R}^{(p_{r})}$, $\tilde{\mathbf{H}}[p_{i}]=\mathbf{H}[p_{r}]$triangleright Replicate embeddings
18:     return $\tilde{\mathbf{H}}$
19:end function
1:number of spectral features $k$ for node embeddings
3:trained model parameters $\theta$
4:function Train($\mathcal{D}$, GNNθ, SignNetθ)
5:     while not converged do
6:         $H\sim\text{Uniform}(\mathcal{D})$triangleright Sample graph
7:         $(B^{(0)},\ldots,B^{(L)})\leftarrow\text{RndRedSeq}(H)$triangleright Sample coarsening sequence by Algorithm 2
8:         $l\sim\text{Uniform}(\{0,\ldots,L\})$triangleright Sample level
9:         if $l=0$ then
10:              $\mathbf{v}_{L}^{(0)}\leftarrow 1$, $\mathbf{v}_{R}^{(0)}\leftarrow 1$
11:         else
13:         end if
14:         if $l=L$ then
16:              $\mathbf{v}_{L}^{(l+1)}\leftarrow 1$
17:              $\mathbf{v}_{R}^{(l+1)}\leftarrow 1$
18:              $\mathbf{e}^{(l)}\leftarrow 1$
19:         end if
21:         $\mathbf{H}^{(l)}\leftarrow\text{Embeddings}(B^{(l+1)},\text{SignNet}_{\theta},\mathbf{v}_{L}^{(l+1)},\mathbf{v}_{R}^{(l+1)})$triangleright Compute node embeddings
25:     end while
26:     return $\theta$
27:end function
1:reduction fraction range $[\rho_{\text{min}},\rho_{\text{max}}]$
2:target hypergraph size $N$, denoising model GNNθ, spectral feature model SignNetθ
3:sampled hypergraph $H=(\mathcal{V},\mathcal{E})$ with $|\mathcal{V}|=N$
4:function Sample($N$, GNNθ, SignNetθ)
5:     $B=(\mathcal{V}_{L},\mathcal{V}_{R},\mathcal{E},\mathbf{f},\mathbf{F}_{L}^{\text{refine}},\mathbf{F}_{R}^{\text{refine}})\leftarrow(\{1\},\{2\},\{(1,2)\},N,0,0)$triangleright Start with a minimal bipartite graph
6:     $\mathbf{v}_{L}\leftarrow[1]$, $\mathbf{v}_{R}\leftarrow[1]$triangleright Initial cluster size vectors
7:     while $|\mathcal{V}_{L}|<N$ do
8:         $\mathbf{H}\leftarrow\text{Embeddings}(B,\text{SignNet}_{\theta},\mathbf{v}_{L},\mathbf{v}_{R})$triangleright Compute node embeddings
9:         $n\leftarrow\|\mathbf{v}_{L}\|_{1}$
10:         $\rho\sim\text{Uniform}([\rho_{\text{min}},\rho_{\text{max}}])$triangleright random reduction fraction
11:         set $n^{+}$ s.t. $n^{+}=\lceil\rho(n+n^{+})\rceil$triangleright number of left side nodes to add
12:         $n^{+}\leftarrow\min(n^{+},N-n)$triangleright ensure not to exceed target size
13:         $\hat{\rho}\leftarrow 1-(n/(n+n^{+}))$triangleright actual reduction fraction
14:         $D_{\theta}\leftarrow\text{GNN}_{\theta}(\cdot,\cdot,\tilde{B}(B,\mathbf{v}_{L},\mathbf{v}_{R}),\mathbf{H},N,\hat{\rho})$
15:         $(\mathbf{v}_{L})_{0},(\mathbf{v}_{R})_{0},(\mathbf{e})_{0},\mathbf{f},\mathbf{F}_{L}^{\text{refine}},\mathbf{F}_{R}^{\text{refine}}\leftarrow\text{Sample}(D_{\theta})$triangleright Sample features
16:         set $\mathbf{v}_{L}$ s.t. for $i\in[n]$: $\mathbf{v}_{L}[i]=2$ if |$|\{j\in[n]\mid(\mathbf{v}_{L})_{0}[j]\geq(\mathbf{v}_{L})_{0}[i]\}|\geq n^{+}$ and $v[i]=1$ otherwise
19:         $B=(\mathcal{V}_{L},\mathcal{V}_{R},\mathcal{E})\leftarrow B(\tilde{B},\mathbf{e},\mathbf{f},\mathbf{F}_{L}^{\text{refine}},\mathbf{F}_{R}^{\text{refine}})$triangleright Refine as per Definition 6
20:     end while
21:     build $H$ from its bipartite representation $B$
22:     return $H$
23:end function
Appendix F Experimental Details
@@
For all experiments, we use embeddings with 32 dimensions for edge selection vectors and node and hyperedge expansion numbers. When they exist, features are embedded with 128 dimensions. SignNet always has 5 layers and a hidden dimension of 128. Positional encodings always have 32 dimensions. We always use 10 layers of Local PPGN. We always use 25 sampling steps. All experiments are run for 24 hours on a single L40S. We use 8 CPU workers.
F.1 Experimental Details: Unfeatured Hypergraphs
@@
Detailed numerical results.
Table 6: Detailed numerical results for ablation studies.
SBM Hypergraphs
77.7% 0.020 0.889 0.041 0.006 5.118 0.015 0.002
79.9% 0.123 0.815 0.023 0.006 6.3 0.015 0.003
71.6% 0.039 1.413 0.005 0.013 9.862 0.006 0.004
81.4% 0.010 0.603 0.005 0.005 3.962 0.009 0.002
Ego Hypergraphs
99.0% 2.383 0.239 0.139 0.006 5.720 0.007 4e-6
98.5% 1.721 0.223 0.184 0.010 3.784 6.7e-4 1.8e-5
100% 0.459 0.129 0.279 0.010 3.010 0 0
100% 0.162 0.171 0.129 0.007 6.061 0 0
Tree Hypergraphs
92.2% 0.005 0.066 0.161 0.011 0.406 0.022 0.006
83.8% 0 0.045 0.083 0.006 0.515 0.025 0.027
74.5% 0.005 0.039 0.050 0.003 0.528 0.042 0.026
82.8% 0 0.043 0.046 0.002 0.258 0.040 0.027
ModelNet Bookshelf
8.615 5.951 0.043 0.027 102.547 0.113 0.003
7.225 4.183 0.109 0.037 63.778 0.197 0.003
1.525 5.232 0.069 0.013 79.016 0.087 0.003
0.875 5.600 0.087 0.019 32.723 0.113 0.002
ModelNet Piano
8.125 3.546 0.075 0.026 42.991 0.102 0.002
2.375 5.529 0.160 0.028 50.349 0.097 0.002
11.275 4.127 0.072 0.027 31.312 0.168 0.002
0.125 1.878 0.028 0.036 43.326 0.072 0.001
F.2 3D Meshes
@@
Metrics. We use the same metrics as for hypergraphs, which are: NodeNumDiff, NodeDegree, EdgeSize, and Spectral. To this we add a metric NearChamDist which computes the Chamfer distance between a point cloud sampled from the surface of the generated mesh and equivalent point clouds sampled from the validation/test set meshes.
@@
Detailed numerical results.
Table 7: Evaluation on Manifold40 Bench and Airplane datasets.
Method Manifold40 Bench Manifold40 Airplane
Cham Dist $\downarrow$ Node Num $\downarrow$ Node Deg $\downarrow$ Edge Size $\downarrow$ Spectral $\downarrow$ Centr Harm $\downarrow$ Centr Clos $\downarrow$ Centr Betw $\downarrow$ Cham Dist $\downarrow$ Node Num $\downarrow$ Node Deg $\downarrow$ Edge Size $\downarrow$ Spectral $\downarrow$ Centr Harm $\downarrow$ Centr Clos $\downarrow$ Centr Betw $\downarrow$
Sequential 0.143 0.367 0.801 0.004 0.007 4.964 0.087 0.006 0.117 0.078 0.332 0.011 0.015 3.131 0.035 0.007

### v2

@@ A.2 Minibatch OT-coupling
$f(\mathbf{x}_{0})=f(\sigma(\mathbf{x}_{0})).$ (14)
@@
Proposition 12 (Shorter paths).
Under the OT-coupling of Algorithm 2, the loss-induced paths for a given minibatch are, on average, shorter than without OT-coupling. Formally, let $\pi$ be the permutation returned by Algorithm 2 for starting points $\mathbf{x}^{0}$ and target points $\mathbf{x}^{1}$. Then
$\sum_{i}\|\mathbf{x}^{0}_{i}-\mathbf{x}^{1}_{i}\|^{2}\;\geq\;\sum_{i}\|\mathbf{x}^{0}_{\pi(i)}-\mathbf{x}^{1}_{i}\|^{2}.$ (19)
Proof.
We can rewrite the left-hand side of (19) as a sum over clusters:
$\sum_{i}\|\mathbf{x}^{0}_{i}-\mathbf{x}^{1}_{i}\|^{2}$ $=\sum_{\text{clusters }\mathcal{C}}\;\sum_{i\in\mathcal{C}}\|\mathbf{x}^{0}_{i}-\mathbf{x}^{1}_{i}\|^{2}.$ (20)
By construction, Algorithm 2 minimizes $\sum_{i\in\mathcal{C}}\|\mathbf{x}^{0}_{i}-\mathbf{x}^{1}_{i}\|^{2}$ within each cluster $\mathcal{C}$.
Therefore,
$\sum_{\text{clusters }\mathcal{C}}\sum_{i\in\mathcal{C}}\|\mathbf{x}^{0}_{i}-\mathbf{x}^{1}_{i}\|^{2}$ $\geq\sum_{\text{clusters }\mathcal{C}}\sum_{i\in\mathcal{C}}\|\mathbf{x}^{0}_{\pi(i)}-\mathbf{x}^{1}_{i}\|^{2}$ (21)
$=\sum_{i}\|\mathbf{x}^{0}_{\pi(i)}-\mathbf{x}^{1}_{i}\|^{2},$ (22)
which proves the claim. ∎
Appendix B Coarsening Sequence Sampling
This section outlines our methodology for sampling a coarsening sequence $\pi\in\Pi_{F}(H)$ for a given hypergraph $H$. The full procedure is detailed in Algorithm 1. At each coarsening step $l$, let $H^{(l)}$ denote the current hypergraph, $B^{(l)}$ its bipartite representation, and $C^{(l)}$ its weighted clique expansion. We begin by sampling a target reduction fraction $\text{red\_frac}\sim\mathcal{U}([\rho_{\text{min}},\rho_{\text{max}}])$. We then evaluate all possible contraction sets $F(C^{(l-1)})$ using a cost function $c$, where lower cost indicates higher preference. We employ a greedy randomized strategy that processes contraction sets in order of increasing cost. For each set:
@@ Appendix B Coarsening Sequence Sampling
i.e., when the number of nodes on the left side (corresponding to the original hypergraph nodes) has been reduced by the sampled fraction. This framework is flexible, allowing a variety of cost functions $c$, contraction families $F$, reduction fraction ranges $\rho_{\text{min}},\rho_{\text{max}}]$, and randomization parameters $\lambda$.
@@
  • Reduction fraction range: $\rho_{\text{min}},\rho_{\text{max}}]=[0.1,0.3]$.
@@
1:contraction family $F$, cost function $c$, reduction fraction range $\rho_{\text{min}},\rho_{\text{max}}]$, randomization parameter $\lambda$
@@
8:  $\mathbf{b}_{L}^{(0)}\leftarrow(1,\dots,1)\in\mathbb{R}^{n}$ $\triangleright$ Initial node budgets
9:  $\mathbf{b}_{R}^{(0)}\leftarrow(1,\dots,1)\in\mathbb{R}^{m}$ $\triangleright$ Initial hyperedge budgets
@@
14:   $\text{red\_frac}\sim\text{Uniform}([\rho_{\text{min}},\rho_{\text{max}}])$ $\triangleright$ Sample reduction fraction
15:   $f\leftarrow c(\cdot,C^{(l-1)},(\mathcal{P}^{(l-1)},\ldots,\mathcal{P}^{(0)}))$ $\triangleright$ Cost function
16:   $\text{accepted\_contractions}\leftarrow\emptyset$
@@
22:        if  $\forall\text{ right cluster }R\in B_{\text{temp}}:|R|\leq 3$ then
23:         $\text{accepted\_contractions}\leftarrow\text{accepted\_contractions}\cup\{S\}$
24:         $C^{(l)}\leftarrow C_{\text{temp}}$, $B^{(l)}\leftarrow B_{\text{temp}}$
@@
26:         $\text{Let }S=\{v_{1},\ldots,v_{k}\},\text{ and the new node be }v^{*}$
@@
Appendix C Examples of Coarsening Sequences
Figure 3: Examples of coarsening sequence for various meshes. Thick lines represent 2-edges.
Appendix D Experimental Details
In this section, we detail all three types of experiments – unfeatured hypergraphs, 3D meshes and graph point clouds – individually, detailing their datasets, baselines, metrics and specific hyperparameters.
@@ Appendix D Experimental Details
For all experiments, we use embeddings with 32 dimensions for edge selection vectors and node and hyperedge expansion numbers. When they exist, features are embedded with 128 dimensions. SignNet always has 5 layers and a hidden dimension of 128. Positional encodings always have 32 dimensions. We always use 10 layers of Local PPGN. We always use 25 sampling steps. All experiments are run for 1M steps on a single L40S. We use 8 CPU workers.
D.1 Unfeatured Hypergraphs
@@
D.2 3D Meshes
@@
Metrics. We use the same metrics as for unfeatured hypergraphs. To this we add a metric ChamDist which computes the minimal Chamfer distance between a point cloud sampled from the surface of the generated mesh and equivalent point clouds sampled from all validation/test set meshes.
@@
D.3 Graph Point Clouds
Datasets. We reuse the same datasets as for 3D meshes. Point clouds comprising 1024 nodes are sampled on the surface of each mesh, then each node is connected to its 3 nearest neighbors. Only the largest connected component is kept.
Metrics. Evaluation Metrics. We adopt the same evaluation protocol as [35] to assess the quality of generated graphs, reporting the following metrics:
  • NumDiff: MMD between node counts distributions of generated and reference graphs.
  • Deg: MMD between degree distributions of generated and reference graphs.
  • Clustering: MMD between clustering coefficient distributions.
  • Orbit: MMD between graphlet orbit count distributions.
  • Spectral: MMD between Laplacian spectra.
  • Wavelet: MMD between wavelet coefficient distributions.
  • Ratio: Average ratio of generated-to-reference values across the above metrics, used as a global indicator of statistical similarity.
To this we add a metric ChamDist which computes the minimal Chamfer distance between the generated mesh and all validation/test set point clouds.
Baselines. We compare against two state-of-the-art flat generative models: Defog [33], DiGress [6]. As these two methods are tailored for discrete data, we use a mixed-discrete continuous framework for continuous features.
Appendix E Implementation Details
E.1 Flow-matching Framework
$\mathcal{L}=\mathbb{E}_{t,\>\mathbf{x}_{0}\sim p_{0},\>\mathbf{x}_{1}\sim p_{1}}\left[\left\|\hat{\mathbf{x}}_{1}(\mathbf{x}_{t},1)-\mathbf{x}_{1}\right\|^{2}\right],$ (25)
where $\mathbf{x}_{t}=(1-t)\mathbf{x}_{0}+t\mathbf{x}_{1}$ is a linear interpolation between samples from the prior and target distributions. Following [37], during training $t$ is sampled using a log-normal distribution, which has been shown to improve performance.
Prior distributions. We use different prior distributions depending on the task:
  • Node and edge predictions: Prior samples $p_{0}$ are drawn from a Gaussian distribution. Targets are either 1-1 or $1$, indicating binary decisions (e.g., whether a node is expanded or an edge is retained).
  • Hyperedge expansion: Similar to node and edge predictions, with targets 1-1, 0, or $1$, encoding the number of expansions (none, one, or two).
  • Budget fractions: Prior samples $p_{0}$ are drawn from a Dirichlet distribution with concentration parameter $\alpha=1.5$, linearly mapped to [1,1][-1,1] via 2x12x-1. The target is the budget fraction of each child node, linearly mapped to [1,1][-1,1] via 2x12x-1. If a cluster is not expanded, the corresponding budget becomes $1$. Parent cluster budgets are encoded using sinusoidal positional encodings [38] (dimension 32, base frequency $10^{-4}$).
  1. i)
  2. ii)
    Find the smallest index $\rho\in\{1,\dots,K\}$ such that:
    $u_{\rho}-\frac{1}{\rho}\left(\sum_{j=1}^{\rho}u_{j}-1\right)>0.$ (27)
  3. iii)
    Compute the threshold:
    $\tau=\frac{1}{\rho}\left(\sum_{j=1}^{\rho}u_{j}-1\right).$ (28)
  4. iv)
    The projection is then:
    $\mathbf{x}^{*}=\max(\mathbf{z}-\tau,0).$ (29)
Graph inpainting. During generation, we use inpainting to enforce constraints: i) budget splits are fixed to 1 for unexpanded clusters or those with a budget of $1$, ii) equal splits are enforced for expanded clusters with a budget of 2, iii) clusters of size one are not allowed to expand; and iv) features of non-expanded clusters are copied unchanged.
E.2 Model Architecture
The denoising network is structured as follows:
  1. 1.
  2. 2.
  3. 3.
    Feature concatenation:
  4. 4.
  5. 5.
    • Edge head: Predicts edge existence.
E.3 Additional Details
Definition 13 (Perturbed hypergraph expansion).
Let $B=(\mathcal{V}_{L},\mathcal{V}_{R},\mathcal{E})$ be a bipartite graph, and let $\mathbf{v}_{L}\in\mathbb{N}^{|\mathcal{V}_{L}|}$ and $\mathbf{v}_{R}\in\mathbb{N}^{|\mathcal{V}_{R}|}$ denote the left and right cluster size vectors. For a given radius $r\in\mathbb{N}$ and probability $0\leq p\leq 1$, we construct $\tilde{B}$ as in Definition 4. Additionally, for each pair of distinct nodes $\mathbf{v}_{L}{(p)}\in\tilde{\mathcal{V}_{L}}$ and $\mathbf{v}_{R}{(q)}\in\tilde{\mathcal{V}}_{R}$ that are within a distance of at most $2r+1$ in $B$, we independently add an edge $e_{\{p_{i},q_{j}\}}$ to $\tilde{\mathcal{E}}$ with probability $p$.
Minibatch OT-coupling. Minibatch OT-coupling [12, 13] accelerates training by jointly sampling priors and targets and reindexing priors to minimize input–output distance, leading to smoother flows and fewer inference steps. As detailed in Section 3.5, we adapt this idea by treating each side of the bipartite graph as a minibatch and applying OT-coupling within them. To preserve the distribution, we restrict permutations to groups of equivalent nodes (identical topology and features, differing only in noise). For efficiency, we assume equivalence only among nodes from the same cluster expansion. Since each cluster produces two or three children, coupling reduces to evaluating two or six possible permutations per cluster, which can be efficiently parallelized with tensors (Algorithm 2).
2:Reindexed noise samples $\{\tilde{z}_{i}\}$
3:Sample noise $\{z_{i}\}$ for each node in $B$
4:for all clusters $V^{(p)}\in B$ do
5:  if $|V^{(p)}|=1$ then
6:   No reassignment needed for singleton cluster
7:  else if $|V^{(p)}|=2$ then
8:   Let $i,j$ be the indices of the two nodes in $V^{(p)}$
9:   Let $x_{i},x_{j}$ be their corresponding targets
10:   Compute normal order cost:
$C_{\text{normal}}\leftarrow\|z_{i}-x_{i}\|^{2}+\|z_{j}-x_{j}\|^{2}$
11:   Compute swapped order cost:
$C_{\text{swap}}\leftarrow\|z_{j}-x_{i}\|^{2}+\|z_{i}-x_{j}\|^{2}$
12:   if $C_{\text{swap}}<C_{\text{normal}}$ then
13:     Swap $z_{i}\leftrightarrow z_{j}$
14:   end if
15:  end if
16:end for
17:return $\{\tilde{z}_{i}\}$ (reassigned noise samples)
Appendix F Training and Sampling Procedures
1:number of spectral features $k$